OpenStudy (anonymous):

If pentagon ABCDE is rotated 180° around the origin to create pentagon A’B’C’D’E’, what is the ordered pair of point C’? (-5, 2) (5, 2) (2, -5) (-2, 5)

OpenStudy (anonymous):

Need the picture of the pentagon or the coordinates of C to answer this one

OpenStudy (anonymous):

Can You Answer a Different Question For me ?

OpenStudy (anonymous):

The bottom shape is a square prism with base edges of 5 feet and a height of 15 feet. What is the volume of the composite figure if the top shape is a right pyramid with a height of 3 feet?

OpenStudy (anonymous):

the total volume will be the sum of the volumes of the square prism and the volume of the pyramid

OpenStudy (anonymous):

25 ft3

OpenStudy (anonymous):

prism volume=5*5*15=375 pyramid volume=(1/3)5*5*3=25 total is 375+25=400 ft^3

OpenStudy (anonymous):

THANKS YOU

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Can You Help me again